OpenStudy (anonymous):

A water tank is h feet high. Water is flowing from this water tank with a velocity, V feet per second. The model representing the relationship between the velocity and height is given by a.) Find the height of a water tank that provides a water flow of 60 feet per second.

OpenStudy (dumbcow):

whats the relationship?

OpenStudy (anonymous):

v=6 sqrt{h}

OpenStudy (dumbcow):

then h = V^2/36 when V = 60 h = 60^2/36 = 100